### v1

# MOC-RVQ: Multilevel Codebook-assisted Digital Generative Semantic Communication

Abstract
Vector quantization-based image semantic communication systems have successfully boosted transmission efficiency, but face a challenge with conflicting requirements between codebook design and digital constellation modulation. Traditional codebooks need a wide index range, while modulation favors few discrete states. To address this, we propose a multilevel generative semantic communication system with a two-stage training framework. In the first stage, we train a high-quality codebook, using a multi-head octonary codebook (MOC) to compress the index range. We also integrate a residual vector quantization (RVQ) mechanism for effective multilevel communication. In the second stage, a noise reduction block (NRB) based on Swin Transformer is introduced, coupled with the multilevel codebook from the first stage, serving as a high-quality semantic knowledge base (SKB) for generative feature restoration. Experimental results highlight MOC-RVQ’s superior performance over methods like BPG or JPEG, even without channel error correction coding.
Index Terms:
vector quantization, generative semantic communication, two-stage training, semantic knowledge base
I Introduction
Traditional communication systems focus on the bit/symbol-level transmission, where the receiver seeks to minimize bit/symbol transmission errors to recover digital media [1]. This allows the design of communication system architecture to be separated from the vast and complex content of media. As a result, bit/symbol-based transmission has long been the prevailing paradigm in traditional communication systems.
Recently, propelled by the rapid development of deep learning technology, a new intelligent transmission paradigm known as semantic communication is gradually gaining prominence. In contrast to traditional transmission, semantic communication aims for the transmission of semantic fidelity [2]. For example, Dai et al. [3] integrate the nonlinear transform as a robust prior to efficiently extract source semantic features. Yang et al. [4] introduce a deep joint source-channel coding (JSCC) strategy designed for the transmission of images in wireless communication. Sun et al. [5] design a multi-level semantic coding and feature transmission mechanism powered by semantic knowledge base. However, these works suffer from efficiency issues, due to the high dimension of latent representation of raw data.
One solution to achieve a compact representation is vector quantization (VQ) technique [6] . Generally, VQ transforms semantic features into a series of indices, providing a more compact format that can be further converted into bits for transmission, then the receiver utilizes the received indices to reconstruct the semantic features through a pre-established learned codebook. Ultimately, the raw data or the task-oriented information can be decoded from these reconstructed semantic features. Nemati et al. [7] delve into the characteristics of VQ-VAE and adapt its training process to formulate a robust JSCC scheme against noisy wireless channels. Fu et al. [8] design a CNN-based transceiver to extract multi-scale semantic features and incorporate multi-scale semantic embedding spaces to facilitate feature quantization. To improve semantic representation in digital transmission under resource limitations, Guo et al. [9] introduce a novel non-linear quantization module with trainable levels for efficient feature extraction.
However, methods above suffer the following two issues:
  • Incompatible issue arises from the disparity between conventional vector quantization and digital constellation modulation. The former, aimed at achieving optimal image representation, employs a learnable codebook with a wide range of code indices. In contrast, the latter is more inclined to handle fewer states (e.g., 16-QAM, 64-QAM, etc.). While it is conceivable to regroup bits representing indices to align with modulation, such an adjustment also disrupts the explicit relationship between indices and constellation points. Note that this explicit relationship is crucial for preserving the local semantic relationship of the underlying neighbour code vectors, which directly affects the quality of reconstructed image.
  • Mismatch issue occurs in the local relationship between code indices and code vectors, which renders vector quantization-based semantic communication systems susceptible to channel noise. For instance, while the difference between index "1" and index "2" is 1, the distance between the underlying code vectors of index "1" and index "2" can be substantial. This mismatch may lead to errors in decoding and, consequently, degrade the overall performance of the communication system.
To tackle the incompatible issue, we introduce MOC-RVQ with two stage training framework. In Stage 1, we introduce (1) MOC, a variant of vector quantizer by constructing a multi-head octonary codebook (MOC) to compress discrete states to a limit of 8, allowing for a direct match with 64-QAM. (2) RVQ, a multi-level semantic transmission mechanism based on residual vector quantization (RVQ), which is helpful to compensate the negative effect of the quantization noise [10, 11]. These two components are integrated to form MOC-RVQ. Subsequently, in Stage 2, we propose a noise reduction block (NRB) and feature requantization to achieve generative feature restoration. For the mismatch issue, we adopt a heuristic codebook reordering algorithm, which is aimed at minimizing the semantic distance between underlying code vectors with nearby indices.
Overall, the main contributions are summarized as follows:
  • We point out two inherent shortcomings of the existing VQ-based semantic communication system and propose a novel two-stage framework.
  • We introduce MOC-RVQ to align with the property of digital constellation modulation. Then noise reduction block (NRB) based on Swin Transformer and feature requantization are incorporated for generative semantic reconstruction.
  • We propose a codebook reordering algorithm, inspired by Gray code mapping, to alleviate the discrepancy in the local relationships among code vectors with nearby indices. This algorithm serves to enhance the overall robustness of the system against channel noise.
II System Model
In this section, we present a general VQ-based semantic communication framework, which we will subsequently extend to our two-stage framework in Section III. The core pipeline is depicted in Fig. 1. In this setup, both the transmitter and receiver employ a pre-established semantic codebook as semantic knowledge base (SKB) to achieve an efficient process for compressing and recovering of semantic features, ultimately leading to a reduction in communication overhead.
II-A Transmitter
Given an input image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, the semantic encoder $E$ (served as understanding knowledge base) is employed initially to transform $\mathbf{I}$ into a semantic feature $\mathbf{z}\in\mathbb{R}^{h\times w\times n_{q}}$. Subsequently, each grid feature $\mathbf{z}^{(i,j)}\in\mathbb{R}^{n_{q}}$ is substituted with the nearest entry in the learnable codebook $\mathcal{C}=\left\{\mathbf{e}_{k}\in\mathbb{R}^{n_{q}}\right\}_{k=0}^{N}$. This substitution yields the quantized feature $\mathbf{z}_{q}$ along with the corresponding code index sequence $\mathbf{s}\in\{0,1,\cdots,N-1\}^{h\cdot w}$, by the following equations:
$\mathbf{z}_{q}^{(i,j)}$ $=\mathbf{Q}\left(\mathbf{z}^{(i,j)}\right)=\mathbf{e}_{\mathbf{s}^{(i,j)}}\in\mathbb{R}^{n_{q}},$ (1)
[…]
where $\mathbf{Q}(\cdot)$ is a quantization operation that maps the continuous encoded grid feature $\mathbf{z}^{(i,j)}$ into the quantized form $\mathbf{z}_{q}^{(i,j)}$ by looking up the closest codebook entry $\mathbf{e}_{k}$. Note that the code index sequence $\mathbf{s}$ can be served as a compact equivalent representation of the quantized feature $\mathbf{z}_{q}$. This implies that transmitting only the code index sequence $\mathbf{s}$ can effectively reduce communication overhead. Following decimal to binary transform and digital constellation modulation, $\mathbf{s}$ is converted as a transmitted complex signal $\mathbf{x}\in\mathbb{C}^{B\times 1}$, where $B$ represents the length of the transmitted signal.
Figure 1: An essential structure for VQ-based semantic communication with the presence of a stochastic physical channel.
Figure 2: Left: The proposed two-stage framework. In Stage 1, we initially pretrain MOC-RVQ and other model components to generate compact representation of high-resolution image. Then the noise reduction block (NRB) is finetuned in Stage 2 to achieve feature restoration. Right: The architecture of the proposed MOC and MOC-RVQ. In MOC, each residual feature $\mathbf{r}_{d}$ is first divided into several heads, then the 8-state codebook is employed to quantize each head feature, and finally these features are concatenated together to form a single quantized feature. In MOC-RVQ, feature are recursive quantized by MOC to produce residual features and the corresponding code indices. Note that the output of MOC-RVQ is a summation over all quantized features $\mathbf{e}^{(d)}_{\mathbf{s}_{d}}$.
II-B Physical Channel
To simplify channel modeling without sacrificing representativeness, we adopt the additive white Gaussian noise (AWGN) as follows:
$\mathbf{y}=\mathbf{x}+\mathbf{w},$ (3)
where $\mathbf{y}\in\mathbb{C}^{B\times 1}$ represents the received complex signal. The AWGN $\mathbf{w}$ has i.i.d elements with zero mean and variance $\sigma^{2}$.
II-C Receiver
After demodulation of the received signal $\mathbf{y}$, the receiver obtains $\hat{\mathbf{x}}$. Subsequently, the binary to decimal transform retrieves the indices $\mathbf{\hat{s}}$ from $\mathbf{\hat{x}}$ (Note that $\mathbf{\hat{s}}$ can be served as a prompt in a large language model, we utilize it to generate high quality reconstructed image). By selecting the corresponding embedding vectors from the codebook $\mathcal{C}$, the receiver reconstructs the semantic feature tensor $\mathbf{\hat{z}}_{q}$, we define this process as semantic feature retrieval $\mathbf{R}(\cdot)$. Finally, leveraging this retrieval semantic information, the semantic decoder $G$ (served as generative knowledge base) generates the reconstructed image $\mathbf{\hat{I}}$.
III The Proposed Method
In this section, we introduce a novel two-stage framework to enhance vector quantization-based semantic communication. Subsequently, we delve into the description of codebook reordering, an algorithm employed to minimize the mismatch between local relationships of code indices and code vectors.
III-A Stage 1: Pretraining of MOC-RVQ
The primary objective of Stage 1 is to train a high-quality encoder/decoder model and codebook without any channel noise, aiming for a compact representation of images. As depicted in Fig. 2 (Left), in Stage 1, an input image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$ is initially fed into the encoder $E$ to yield a continuous semantic feature $\mathbf{z}\in\mathbb{R}^{h\times w\times n_{q}}$. Subsequently, the MOC-RVQ layer quantizes this feature, resulting in the quantized feature $\mathbf{z}_{q}$ and an associated code indices sequence $\mathbf{s}\in\{0,\cdots,N-1\}^{h\cdot w\cdot P\cdot D}$. Here, $P$ and $D$ represent the number of heads and the level of quantizations respectively, and $N$ is set to 8 to align with the few states prior of 64-QAM. Finally, the quantized feature $\mathbf{z}_{q}$ is fed into the decoder $G$ to reconstruct the image as $\mathbf{\hat{I}}$.
The detailed designs of multi-head octonary codebook (MOC) and MOC-RVQ are illustrated in Fig. 2 (Right), and their functionality can be described as follows:
MOC
Given a quantization level $d$, each residual feature $\mathbf{r}_{d}\in\mathbb{R}^{h\times h\times n_{q}}$ is initially divided into several heads $\mathbf{r}_{d,i}\in\mathbb{R}^{h\times h\times n_{q}/P}$ (typically set $P$ to 4 by default). Subsequently, an 8-state codebook is employed to quantize each head, yielding head-wise quantized features $\mathbf{e}_{d,i}$. These quantized features are then concatenated to form a single quantized feature $\mathbf{e}_{s_{d}}^{(d)}$. This distinctive codebook design for digital constellation modulation is termed as the multi-head octonary codebook (MOC). In contrast to conventional codebook with size $K$, our proposed MOC dramatically expands the feature matching space, generating $N^{P}$ discrete states through the combination of each head state.
MOC-RVQ
By replacing the naive quantizer with MOC, we introduce a multilevel semantic transmission mechanism based on residual vector quantization (RVQ) to counteract potential adverse effects resulting from the quantization process, ultimately further enhances the system’s reconstruction performance. Specifically, with a default quantization level $D$ set to 4, the semantic feature $z$ undergoes recursive quantization to produce residual code indices as follows:
$\mathcal{MOC-RVQ}(\mathbf{z})=\left(\mathbf{s}_{1},\cdots,\mathbf{s}_{D}\right),$ (4)
where $\mathbf{s}_{d}\in\{0,\cdots,N-1\}^{h\cdot w\cdot P}$ represents the code indices of $\mathbf{z}$ at level $d$, and $P$ denotes the number of heads. Given an initial residual $\mathbf{r}_{0}=\mathbf{z}$, the recursive process computes indices $s_{d}$ and the next residual $\mathbf{r}_{d}$ as follows:
$\mathbf{s}_{d}$ $=\mathcal{MOC}^{(d)}(\mathbf{r}_{d-1}),$ (5)
$\mathbf{r}_{d}$ $=\mathbf{r}_{d-1}-\mathbf{e}^{(d)}_{\mathbf{s}_{d}},$ (6)
where $\mathcal{MOC}^{(d)}(\cdot)$ is the $d$-th MOC quantizer, and $\mathbf{e}^{(d)}$ is the corresponding codebook embedding. The final quantized result of MOC-RVQ is obtained by $\mathbf{z}_{q}=\sum_{d=1}^{D}\mathbf{e}^{(d)}_{\mathbf{s}_{d}}$.
Training Objectives
To train the proposed model, several loss functions are introduced into our method. Since the feature quantization operation is non-differentiable, we adopt the approach from  [12] by copying gradients from $G$ to $E$ during backpropagation. This enables end-to-end training of the encoder/decoder model and codebook, guided by the following objective function:
$\mathcal{L}^{\prime}_{VQ}(E,G,\mathcal{C})=\left\|\mathbf{\hat{I}}-\mathbf{I}\right\|_{1}+\sum_{d,h}\|\operatorname{sg}[\mathbf{r}_{d,h}]-\mathbf{e}_{d,h}\|_{2}^{2},$ (7)
where $sg[\cdot]$ represents the stop-gradient operation, and $\mathbf{r}_{d,h}$ and $\mathbf{e}_{d,h}$ denote the d-th quantization of the h-th head of residual and quantized features, respectively.
To enhance texture restoration, a semantic guidance loss, as proposed in  [13], is also introduced as a regularization term:
$\mathcal{L}_{VQ}=\mathcal{L}_{VQ}^{\prime}+\gamma\|\operatorname{CONV}(\mathbf{z})-\phi(\mathbf{I})\|_{2}^{2},$ (8)
where $CONV(\cdot)$ denotes a $1\times 1$ convolution layer to align the dimensions of $\mathbf{z}$, $\phi$ represents the pretrained VGG19 model, and $\gamma$ is set to 0.1. Additionally, in accordance with [12], perceptual loss and adversarial loss are employed for model pretraining.
III-B Stage 2: Finetuning for Noise Reduction
Since Stage 1 is trained end-to-end without accounting for channel noise, leading to suboptimal reconstructed results in the presence of noise, we integrate a noise reduction block (NRB) based on a stack of residual Swin Transformer layers, as mentioned in [13] into decoder $G$ to conduct generative image reconstruction.
Building upon the pretrained model and codebook from Stage 1, as depicted in Fig. 2 (Left), Stage 2 introduces an physical communication channel module (refer to Section II-B for details) between transmitter and receiver for physical channel simulation. Then the noise reduction block is incorporated after the retrieval feature $\mathbf{\hat{z}}_{q}$ to refine semantic features, yielding the refined semantic feature $\mathbf{\hat{z}}_{r}$.
Feature Requantization
To further enhance the quality of $\mathbf{\hat{z}}_{r}$, we propose requantizing it to obtain $\mathbf{\tilde{z}}_{r}$, formulated as follows:
$\mathbf{\tilde{z}}_{r}=\mathcal{MOC-RVQ}(\mathbf{\hat{z}}_{r}).$ (9)
The intuition behind this lies in the fact that MOC is trained in the absence of any noise, establishing itself as a high-quality semantic knowledge base to inject high quality prior for better feature restoration.
Training Objectives
For NRB finetuning, we generate the noiseless quantized semantic feature of image $\mathbf{I}$ through $\mathbf{z}_{q}=\mathcal{MOC}(E(\mathbf{I}))$ as the ground-truth feature. This ground-truth feature is then used to calculate the Gram matrix loss and L2 loss for the refined semantic feature $\mathbf{\hat{z}}_{r}$ by the following formula:
$\mathcal{L}_{NR}=\|\psi(\mathbf{\hat{z}}_{r})-\psi(sg[\mathbf{z}_{q}])\|_{2}^{2}+\alpha\|\mathbf{\hat{z}}_{r}-sg[\mathbf{z}_{q}]\|_{2}^{2},$ (10)
where $\psi(\cdot)$ is the flattened version of the Gram matrix of a given feature, and $\alpha$ is set to $0.25$. The first term of $\mathcal{L}_{NR}$, also known as style loss, has been demonstrated to be effective for improved feature recovery [13]. Note that the model and codebook trained in Stage 1 remain fixed, and we only finetune the NRB for noise reduction.
III-C Codebook Reordering
As discussed in Section I, a discrepancy exists in the local relationship between code indices and code vectors, rendering VQ-based semantic communication systems susceptible to channel noise. In response to this challenge, we propose a heuristic solution through the codebook reordering (CR) algorithm outlined in Algorithm 1. The algorithm systematically searches for the nearest neighbor vector within $\mathcal{C}$, constructing a new vector sequence, which is then assembled to form a preliminarily sorted codebook. Furthermore, to align with the design principles of Gray code mapping in digital constellation modulation, we also integrate Gray code mapping into our codebook reordering algorithm. The basic idea is to establish proximity relationships under the context of Gray code mapping. For instance, given a 3-bit Gray code sequence represented in decimals as $g=\{0,1,3,2,6,7,5,4\}$, it becomes apparent that "2" and "6" are neighboring elements within the framework of Gray code mapping. Thus, it is essential to allocate two closely related code vectors to these indices.
input : A disordered codebook $\mathcal{C}$ of size $N\times n_{q}$
output : An ordered codebook $\mathcal{C}^{*}$ of size $N\times n_{q}$
// Initialization
1 $\mathcal{C}^{*}\leftarrow\varnothing$ ;
2 $\mathbf{s}_{0}\leftarrow$ mean of $\mathcal{C}$ along dimension 0 ;
3 for $i\leftarrow 1$ to $N$ do
4       $\mathbf{s}_{i}\leftarrow$ the vector closest to $\mathbf{s}_{i-1}$ in $\mathcal{C}$ ;
5       Remove $\mathbf{s}_{i}$ from $\mathcal{C}$ ;
6       Append $\mathbf{s}_{i}$ to $\mathcal{C}^{*}$ ;
       // Gray Code Mapping
7       $g\leftarrow$ generate Gray code sequence of size $N$ ;
8       $\mathcal{C}^{*}[g,:]\leftarrow$ $\mathcal{C}^{*}$ ;
9
Algorithm 1 Codebook Reordering
After codebook reordering, the semantic distance between two adjacent Gray code vectors in the updated codebook $\mathcal{C}^{*}$ is reduced compared to the previous disordered version. This addresses the issue where a small jump in the code index could lead to a significant jump in the semantic vector.
IV Experiments
(a) PSNR (b) SSIM (c) LPIPS
Figure 3: Experimental comparison of methods using PSNR, SSIM, and LPIPS metrics over an additive white Gaussian noise (AWGN) channel with SNR varying from -5 to 30. L1 to L4 denote the level of quantization of our proposed MOC-RVQ, which directly determines the number of bits that need to be transmitted for communication.
(a) PSNR (b) SSIM (c) LPIPS
Figure 4: Ablation study to comprehend the influence of our proposed noise reduction block (NRB) and codebook reordering (CR) algorithm. Note that all experiments are conducted on the basis of L4 transmission.
IV-A Implementation Details
Model and Dataset
In accordance with Chen et al. [13], we employ a baseline autoencoder with a downsampling factor $f=8$, indicating that the resolution of latent features is $1/f^{2}$ of the input resolution. To ensure effective model training, we assemble a comprehensive training set by integrating diverse datasets: DIV2K [14], Flickr2K [15], DIV8K [16], and a dataset comprising 10,000 facial images from FFHQ [17]. The generation of training patches involves the following steps: (1) extraction of non-overlapping patches with dimensions $512\times 512$; (2) filtration of patches with minimal texture variation; (3) for the FFHQ dataset, a random resizing operation is applied with scale factors in the range [0.5, 1.0] before cropping. For testing, we uniformly sample 25 high resolution images (with average resolution of $2012\times 1407$) from DIV2K validation set as our test set to simulate modern image transmission scenarios.
Training Details
We employ the Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.99$ for both stages, maintaining a learning rate of $0.0001$ throughout the training process. The input image size is randomly cropped to $256\times 256$ with a batch size of 16. The model pretraining stage requires approximately 3 days on 2 GeForce RTX 3090 GPUs, and the NRB finetuning stage also takes about 3 days on the same device.
IV-B Simulations and Discussions
Figure 5: Visualizing the impact of NRB across varying channel conditions. Even in poor channel quality (SNR=5), our full model impressively reconstructs images with clear semantic meaning, demonstrating the robustness of VQ-based semantic communication systems with the assistance of NRB.
In this section, we empirically demonstrate the qualitative and quantitative results of our approach, conduct model ablations, and compare our method with two state-of-the-art baselines. Note that all approaches are under 64-QAM. For channel simulation, initially, we simulate an additive white Gaussian noise (AWGN) channel with SNR varying from -5 to 30 to encompass low, medium, and high-quality channel conditions. Subsequently, we finetune our proposed NRB under this AWGN channel to enhance image reconstruction.
For comparison, we apply the better portable graphics (BPG [18]) and the joint picture expert group (JPEG [19]) as baselines. Notably, both of these image codec formats are susceptible to bit errors, leading to potential image format errors and decoding failures. To mitigate this, we introduce a widely used channel error correction code called Low-Density Parity-Check (LDPC) by setting the LDPC block length and bit rate as 648 bits and 1/2, respectively. Thanks to the self-restoring capability of our proposed method, we do not apply any channel error correction code, which reduces the number of transmitted bits and significantly improves the transmission efficiency of the proposed communication system.
As depicted in Fig. 3, we utilize PSNR, SSIM, and LPIPS as metrics for various methods with a similar number of transmitted symbols over the AWGN channel. LPIPS [20] is specifically introduced for evaluating the perceptual quality of generated images in semantic communication scenarios. Our proposed method exhibits remarkable robustness in deteriorating channel quality, outperforming traditional image coding even without channel error correction mechanism. This is attributed to the VQ-based semantic communication system’s independence from encoding formats, where minor bit changes do not lead to decoding failures. Additionally, the incorporation of NRB and codebook reordering further enhances the robustness of the semantic communication system.
Figure 3 also reveals that lowering the transmission level (L4 to L1) of our model reduces the number of transmitted bits, but it also compromises reconstruction performance. Especially, we can clearly observe the gaps at different levels in LPIPS metrics, which implies that we can dynamically adjust the number of bits that need to be transmitted based on the quality of the channel and the receiver’s requirements for image quality.
In Figure 4, we present an ablation study to understand the impact of our proposed NRB and the CR algorithm. Focusing on LPIPS scores, which better capture perceptual quality, we observe the crucial role of NRB in reconstruction performance, especially under poor channel quality. Simultaneously, the proposed codebook reordering algorithm provides additional gains in image reconstruction performance.
Figure 5 illustrates the visual effect of NRB. With the assistance of NRB, even under relatively poor channel quality (SNR=5), our full model impressively reconstructs images with clear semantic meaning, which greatly expands the working conditions of VQ-based semantic communication systems.
V Conclusion
We introduce MOC-RVQ, a novel two-stage framework addressing challenges in vector quantization and digital constellation modulation for semantic communication systems. By incorporating a multi-head octonary codebook (MOC) to compress the index range and a residual vector quantization (RVQ) mechanism for efficient multilevel semantic communication, our approach demonstrates superior performance over existing methods like BPG and JPEG, even without channel error correction coding. The innovative two-stage model training framework, featuring a noise reduction block (NRB), enhances the system’s ability to restore high-quality semantic features in the presence of channel noise. Additionally, a heuristic codebook reordering algorithm improves system robustness against channel noise, mitigating the inherent mismatch between vector quantization and digital constellation modulation. MOC-RVQ represents a significant advancement in generative semantic communication, providing a promising solution for the joint design challenges and paving the way for further exploration in intelligent communication systems.

### v2

@@
 Abstract
-Vector quantization-based image semantic communication systems have successfully boosted transmission efficiency, but face a challenge with conflicting requirements between codebook design and digital constellation modulation.
-Traditional codebooks need a wide index range, while modulation favors few discrete states.
+Vector quantization-based image semantic communication systems have successfully boosted transmission efficiency, but face challenges with conflicting requirements between codebook design and digital constellation modulation.
+Traditional codebooks need wide index ranges, while modulation favors few discrete states.
 To address this, we propose a multilevel generative semantic communication system with a two-stage training framework.
 In the first stage, we train a high-quality codebook, using a multi-head octonary codebook (MOC) to compress the index range.
-We also integrate a residual vector quantization (RVQ) mechanism for effective multilevel communication.
+In addition, a residual vector quantization (RVQ) mechanism is also integrated for effective multilevel communication.
 In the second stage, a noise reduction block (NRB) based on Swin Transformer is introduced, coupled with the multilevel codebook from the first stage, serving as a high-quality semantic knowledge base (SKB) for generative feature restoration.
-Experimental results highlight MOC-RVQ’s superior performance over methods like BPG or JPEG, even without channel error correction coding.
+Finally, to simulate modern image transmission scenarios, we employ a diverse collection of high-resolution 2K images as the test set.
+The experimental results consistently demonstrate the superior performance of MOC-RVQ over conventional methods such as BPG or JPEG.
+Additionally, MOC-RVQ achieves comparable performance to an analog JSCC scheme, while needing only one-sixth of the channel bandwidth ratio (CBR) and being directly compatible with digital transmission systems.
 Index Terms:
-vector quantization, generative semantic communication, two-stage training, semantic knowledge base
+Vector Quantization; Generative Semantic Communication; Semantic Knowledge Base; Image Transmission
 I Introduction
-Traditional communication systems focus on the bit/symbol-level transmission, where the receiver seeks to minimize bit/symbol transmission errors to recover digital media [1].
-This allows the design of communication system architecture to be separated from the vast and complex content of media.
-As a result, bit/symbol-based transmission has long been the prevailing paradigm in traditional communication systems.
-Recently, propelled by the rapid development of deep learning technology, a new intelligent transmission paradigm known as semantic communication is gradually gaining prominence.
+Propelled by the rapid development of deep learning technology, a new intelligent transmission paradigm known as semantic communication is gradually gaining prominence.
 In contrast to traditional transmission, semantic communication aims for the transmission of semantic fidelity [1].
 For example, Dai et al.
-[3] integrate the nonlinear transform as a robust prior to efficiently extract source semantic features.
-Yang et al.
-[4] introduce a deep joint source-channel coding (JSCC) strategy designed for the transmission of images in wireless communication.
+[2] integrate the nonlinear transform as a robust prior to efficiently extract source semantic features for image transmission.
 Sun et al.
 [3] design a multi-level semantic coding and feature transmission mechanism powered by semantic knowledge base.
+Wu et al.
+[4] propose channel denoising diffusion models (CDDM) for wireless communications by developing specific training and sampling algorithms tailored to the forward diffusion process and the reverse sampling process.
 However, these works suffer from efficiency issues, due to the high dimension of latent representation of raw data.
-One solution to achieve a compact representation is vector quantization (VQ) technique [6] .
-Generally, VQ transforms semantic features into a series of indices, providing a more compact format that can be further converted into bits for transmission, then the receiver utilizes the received indices to reconstruct the semantic features through a pre-established learned codebook.
+One solution to achieve compact representation is vector quantization (VQ) technique [5] .
+Generally, VQ transforms semantic features into a series of indices, providing a more compact format that can be further converted into bits for transmission, then the receiver utilizes the received indices to reconstruct the semantic features through a pre-trained semantic codebook.
 Ultimately, the raw data or the task-oriented information can be decoded from these reconstructed semantic features.
-Nemati et al.
+For instance, Nemati et al.
 [6] delve into the characteristics of VQ-VAE and adapt its training process to formulate a robust JSCC scheme against noisy wireless channels.
@@
 [7] design a CNN-based transceiver to extract multi-scale semantic features and incorporate multi-scale semantic embedding spaces to facilitate feature quantization.
-To improve semantic representation in digital transmission under resource limitations, Guo et al.
-[9] introduce a novel non-linear quantization module with trainable levels for efficient feature extraction.
-However, methods above suffer the following two issues:
+However, the VQ methods mentioned above suffer the following two issues:
   • Incompatible issue arises from the disparity between conventional vector quantization and digital constellation modulation.
-The former, aimed at achieving optimal image representation, employs a learnable codebook with a wide range of code indices.
+The former, aimed at achieving optimal image representation, normally employs a learnable codebook with wide index ranges.
 In contrast, the latter is more inclined to handle fewer states (e.g., 16-QAM, 64-QAM, etc.).
 While it is conceivable to regroup bits representing indices to align with modulation, such an adjustment also disrupts the explicit relationship between indices and constellation points.
-Note that this explicit relationship is crucial for preserving the local semantic relationship of the underlying neighbour code vectors, which directly affects the quality of reconstructed image.
+Note that this explicit correspondence is crucial for preserving the local semantic relationship of the underlying neighbour code vectors, which directly affects the quality of reconstructed image.
   • Mismatch issue occurs in the local relationship between code indices and code vectors, which renders vector quantization-based semantic communication systems susceptible to channel noise.
@@
 This mismatch may lead to errors in decoding and, consequently, degrade the overall performance of the communication system.
-To tackle the incompatible issue, we introduce MOC-RVQ with two stage training framework.
-In Stage 1, we introduce (1) MOC, a variant of vector quantizer by constructing a multi-head octonary codebook (MOC) to compress discrete states to a limit of 8, allowing for a direct match with 64-QAM.
+To tackle the incompatible issue, we introduce MOC-RVQ with a two-stage training framework.
+In Stage 1, we introduce: (1) MOC, a variant of vector quantizer by constructing a multi-head octonary codebook (MOC) to compress the range of indices to a limit of 8, allowing for a direct match with 64-QAM.
 (2) RVQ, a multi-level semantic transmission mechanism based on residual vector quantization (RVQ), which is helpful to compensate the negative effect of the quantization noise [8, 9].
-These two components are integrated to form MOC-RVQ.
-Subsequently, in Stage 2, we propose a noise reduction block (NRB) and feature requantization to achieve generative feature restoration.
+We combine MOC and RVQ to form MOC-RVQ.
+In Stage 2, we introduce a noise reduction block (NRB) along with feature requantization to enable generative feature restoration.
 For the mismatch issue, we adopt a heuristic codebook reordering algorithm, which is aimed at minimizing the semantic distance between underlying code vectors with nearby indices.
 Overall, the main contributions are summarized as follows:
-  • We point out two inherent shortcomings of the existing VQ-based semantic communication system and propose a novel two-stage framework.
-  • We introduce MOC-RVQ to align with the property of digital constellation modulation.
-Then noise reduction block (NRB) based on Swin Transformer and feature requantization are incorporated for generative semantic reconstruction.
-  • We propose a codebook reordering algorithm, inspired by Gray code mapping, to alleviate the discrepancy in the local relationships among code vectors with nearby indices.
+  • We point out two inherent shortcomings of the existing VQ-based semantic communication systems and propose a novel two-stage training framework.
+  • We design MOC-RVQ to align with the preference of digital constellation modulation.
+Then the noise reduction block (NRB) based on Swin Transformer and feature requantization are incorporated for generative semantic reconstruction.
+  • We develop a codebook reordering algorithm to mitigate inconsistencies in the local relationships among code vectors with adjacent indices.
 This algorithm serves to enhance the overall robustness of the system against channel noise.
-II System Model
-In this section, we present a general VQ-based semantic communication framework, which we will subsequently extend to our two-stage framework in Section III.
-The core pipeline is depicted in Fig.
+II Preliminary
+In this section, we present a general VQ-based semantic communication framework as preliminary.
+Note that this framework will be expanded and adjusted into our proposed two-stage framework in Section III.
+The main process is illustrated in Fig.
 1.
-In this setup, both the transmitter and receiver employ a pre-established semantic codebook as semantic knowledge base (SKB) to achieve an efficient process for compressing and recovering of semantic features, ultimately leading to a reduction in communication overhead.
+In this design, both the transmitter and receiver utilize a shared pre-trained semantic codebook to efficiently compress and reconstruct semantic features, thereby reducing communication overhead.
 II-A Transmitter
-Given an input image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, the semantic encoder $E$ (served as understanding knowledge base) is employed initially to transform $\mathbf{I}$ into a semantic feature $\mathbf{z}\in\mathbb{R}^{h\times w\times n_{q}}$.
-Subsequently, each grid feature $\mathbf{z}^{(i,j)}\in\mathbb{R}^{n_{q}}$ is substituted with the nearest entry in the learnable codebook $\mathcal{C}=\left\{\mathbf{e}_{k}\in\mathbb{R}^{n_{q}}\right\}_{k=0}^{N}$.
+Given an input image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, the semantic encoder $E$ (serves as understanding knowledge base) is employed initially to transform $\mathbf{I}$ into semantic feature $\mathbf{z}\in\mathbb{R}^{h\times w\times n_{q}}$.
+Subsequently, each grid feature $\mathbf{z}^{(i,j)}\in\mathbb{R}^{n_{q}}$ is substituted with the nearest vector in the learnable codebook $\mathcal{C}=\left\{\mathbf{e}_{k}\in\mathbb{R}^{n_{q}}\right\}_{k=0}^{N}$.
 This substitution yields the quantized feature $\mathbf{z}_{q}$ along with the corresponding code index sequence $\mathbf{s}\in\{0,1,\cdots,N-1\}^{h\cdot w}$, by the following equations:
+[…]
 $\mathbf{z}_{q}^{(i,j)}$ $=\mathbf{Q}\left(\mathbf{z}^{(i,j)}\right)=\mathbf{e}_{\mathbf{s}^{(i,j)}}\in\mathbb{R}^{n_{q}},$ (2)
-[…]
-where $\mathbf{Q}(\cdot)$ is a quantization operation that maps the continuous encoded grid feature $\mathbf{z}^{(i,j)}$ into the quantized form $\mathbf{z}_{q}^{(i,j)}$ by looking up the closest codebook entry $\mathbf{e}_{k}$.
-Note that the code index sequence $\mathbf{s}$ can be served as a compact equivalent representation of the quantized feature $\mathbf{z}_{q}$.
-This implies that transmitting only the code index sequence $\mathbf{s}$ can effectively reduce communication overhead.
-Following decimal to binary transform and digital constellation modulation, $\mathbf{s}$ is converted as a transmitted complex signal $\mathbf{x}\in\mathbb{C}^{B\times 1}$, where $B$ represents the length of the transmitted signal.
+where, $\mathbf{Q}(\cdot)$ represents a quantization operation that maps the continuous encoded grid feature $\mathbf{z}^{(i,j)}$ to its quantized form $\mathbf{z}_{q}^{(i,j)}$ by selecting the closest codebook vector $\mathbf{e}_{k}$.
+Notably, the code index sequence $\mathbf{s}$ can serve as a compact equivalent representation of the quantized feature $\mathbf{z}_{q}$, meaning that transmitting only the code index sequence $\mathbf{s}$ can effectively reduce communication overhead.
+After applying decimal-to-binary transformation and digital constellation modulation, $\mathbf{s}$ is transformed into a transmitted complex signal $\mathbf{x}\in\mathbb{C}^{B\times 1}$, where $B$ denotes the length of the transmitted signal.
 Figure 1: An essential structure for VQ-based semantic communication with the presence of a stochastic physical channel.
@@
 Right: The architecture of the proposed MOC and MOC-RVQ.
-In MOC, each residual feature $\mathbf{r}_{d}$ is first divided into several heads, then the 8-state codebook is employed to quantize each head feature, and finally these features are concatenated together to form a single quantized feature.
-In MOC-RVQ, feature are recursive quantized by MOC to produce residual features and the corresponding code indices.
+In MOC, each residual feature $\mathbf{r}_{d-1}$ is first divided into several heads, then the 8-state codebooks are employed to quantize each head feature, and finally these features are concatenated together to form a single quantized feature.
+In MOC-RVQ, feature are recursive quantized by MOCs to produce residual features and the corresponding code indices.
 Note that the output of MOC-RVQ is a summation over all quantized features $\mathbf{e}^{(d)}_{\mathbf{s}_{d}}$.
@@
 II-C Receiver
-After demodulation of the received signal $\mathbf{y}$, the receiver obtains $\hat{\mathbf{x}}$.
-Subsequently, the binary to decimal transform retrieves the indices $\mathbf{\hat{s}}$ from $\mathbf{\hat{x}}$ (Note that $\mathbf{\hat{s}}$ can be served as a prompt in a large language model, we utilize it to generate high quality reconstructed image).
-By selecting the corresponding embedding vectors from the codebook $\mathcal{C}$, the receiver reconstructs the semantic feature tensor $\mathbf{\hat{z}}_{q}$, we define this process as semantic feature retrieval $\mathbf{R}(\cdot)$.
-Finally, leveraging this retrieval semantic information, the semantic decoder $G$ (served as generative knowledge base) generates the reconstructed image $\mathbf{\hat{I}}$.
+Following the demodulation of the received signal $\mathbf{y}$ and the subsequent binary-to-decimal transformation, the receiver retrieves the indices $\mathbf{\hat{s}}$ from $\mathbf{y}$.
+It is noteworthy that $\mathbf{\hat{s}}$ can be analogized to the prompt in large language models, and we employ it to generate high-quality reconstructed images.
+Specifically, by selecting the corresponding embedding vectors from the codebook $\mathcal{C}$ using the received indices $\mathbf{\hat{s}}$, the receiver reconstructs the semantic feature $\mathbf{\hat{z}}_{q}$, we define this process as semantic feature retrieval $\mathbf{R}(\cdot)$.
+Then, leveraging the retrieval semantic feature $\mathbf{\hat{z}}_{q}$, the semantic decoder $G$ (serves as generative knowledge base) generates the reconstructed image $\mathbf{\hat{I}}$.
 III The Proposed Method
-In this section, we introduce a novel two-stage framework to enhance vector quantization-based semantic communication.
+In this section, we expand the pipeline mentioned in Section II and introduce a novel two-stage training framework to enhance VQ-based semantic communication.
 Subsequently, we delve into the description of codebook reordering, an algorithm employed to minimize the mismatch between local relationships of code indices and code vectors.
 III-A Stage 1: Pretraining of MOC-RVQ
-The primary objective of Stage 1 is to train a high-quality encoder/decoder model and codebook without any channel noise, aiming for a compact representation of images.
+The primary objective of Stage 1 is to train a high-quality encoder/decoder model and multilevel codebook without any channel noise, aiming for a compact representation of image.
 As depicted in Fig.
 2 (Left), in Stage 1, an input image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$ is initially fed into the encoder $E$ to yield a continuous semantic feature $\mathbf{z}\in\mathbb{R}^{h\times w\times n_{q}}$.
-Subsequently, the MOC-RVQ layer quantizes this feature, resulting in the quantized feature $\mathbf{z}_{q}$ and an associated code indices sequence $\mathbf{s}\in\{0,\cdots,N-1\}^{h\cdot w\cdot P\cdot D}$.
-Here, $P$ and $D$ represent the number of heads and the level of quantizations respectively, and $N$ is set to 8 to align with the few states prior of 64-QAM.
+Subsequently, the MOC-RVQ layer quantizes this feature, resulting in the quantized feature $\mathbf{z}_{q}$ and the associated code indices sequence $\mathbf{s}\in\{0,\cdots,N-1\}^{h\cdot w\cdot P\cdot D}$.
+Here, $P$ and $D$ represent the number of heads and the level of quantizations respectively, and $N$ is set to 8 to directly align with the few states preference of 64-QAM.
 Finally, the quantized feature $\mathbf{z}_{q}$ is fed into the decoder $G$ to reconstruct the image as $\mathbf{\hat{I}}$.
@@
 MOC
-Given a quantization level $d$, each residual feature $\mathbf{r}_{d}\in\mathbb{R}^{h\times h\times n_{q}}$ is initially divided into several heads $\mathbf{r}_{d,i}\in\mathbb{R}^{h\times h\times n_{q}/P}$ (typically set $P$ to 4 by default).
-Subsequently, an 8-state codebook is employed to quantize each head, yielding head-wise quantized features $\mathbf{e}_{d,i}$.
+Given a quantization level $d$, each residual feature $\mathbf{r}_{d-1}\in\mathbb{R}^{h\times w\times n_{q}}$ is initially divided into several heads $\mathbf{r}_{d-1,i}\in\mathbb{R}^{h\times w\times n_{q}/P}$ (typically set $P$ to 4 by default).
+Subsequently, the 8-state codebooks are employed to quantize each head separately, resulting in head-wise quantized features $\mathbf{e}_{d,i}$.
 These quantized features are then concatenated to form a single quantized feature $\mathbf{e}_{s_{d}}^{(d)}$.
-This distinctive codebook design for digital constellation modulation is termed as the multi-head octonary codebook (MOC).
-In contrast to conventional codebook with size $K$, our proposed MOC dramatically expands the feature matching space, generating $N^{P}$ discrete states through the combination of each head state.
+We name this distinctive codebook design for aliging the preference of digital constellation modulation as multi-head octonary codebook (MOC).
+It is worth noting that, unlike a conventional codebook of size $K$, our proposed MOC significantly expands the feature matching space by generating $N^{P}$ discrete states through the combination of individual head states.
 MOC-RVQ
@@
 Specifically, with a default quantization level $D$ set to 4, the semantic feature $z$ undergoes recursive quantization to produce residual code indices as follows:
-$\mathcal{MOC-RVQ}(\mathbf{z})=\left(\mathbf{s}_{1},\cdots,\mathbf{s}_{D}\right),$ (4)
+$\mathcal{MOC-RVQ}(\mathbf{z})=\left(\mathbf{s}_{1},\cdots,\mathbf{s}_{d},\cdots,\mathbf{s}_{D}\right),$ (4)
 where $\mathbf{s}_{d}\in\{0,\cdots,N-1\}^{h\cdot w\cdot P}$ represents the code indices of $\mathbf{z}$ at level $d$, and $P$ denotes the number of heads.
-Given an initial residual $\mathbf{r}_{0}=\mathbf{z}$, the recursive process computes indices $s_{d}$ and the next residual $\mathbf{r}_{d}$ as follows:
+More specifically, given an initial residual $\mathbf{r}_{0}=\mathbf{z}$, the recursive process computes indices $s_{d}$ and the next residual $\mathbf{r}_{d}$ as follows:
 $\mathbf{s}_{d}$ $=\mathcal{MOC}^{(d)}(\mathbf{r}_{d-1}),$ (5)
@@
 This enables end-to-end training of the encoder/decoder model and codebook, guided by the following objective function:
-$\mathcal{L}^{\prime}_{VQ}(E,G,\mathcal{C})=\left\|\mathbf{\hat{I}}-\mathbf{I}\right\|_{1}+\sum_{d,h}\|\operatorname{sg}[\mathbf{r}_{d,h}]-\mathbf{e}_{d,h}\|_{2}^{2},$ (7)
-where $sg[\cdot]$ represents the stop-gradient operation, and $\mathbf{r}_{d,h}$ and $\mathbf{e}_{d,h}$ denote the d-th quantization of the h-th head of residual and quantized features, respectively.
+$\mathcal{L}^{\prime}_{VQ}(E,G,\mathcal{C})=\left\|\mathbf{\hat{I}}-\mathbf{I}\right\|_{1}+\sum_{d,h}\|\operatorname{sg}[\mathbf{r}_{d-1,h}]-\mathbf{e}_{d,h}\|_{2}^{2},$ (7)
+where $sg[\cdot]$ represents the stop-gradient operation.
+Note that the first term of $\mathcal{L}^{\prime}_{VQ}$ stands for the image-level $L_{1}$ loss, while the second term is employed for updating the proposed multilevel octonary codebooks.
 To enhance texture restoration, a semantic guidance loss, as proposed in  [11], is also introduced as a regularization term:
@@
 where $CONV(\cdot)$ denotes a $1\times 1$ convolution layer to align the dimensions of $\mathbf{z}$, $\phi$ represents the pretrained VGG19 model, and $\gamma$ is set to 0.1.
-Additionally, in accordance with [12], perceptual loss and adversarial loss are employed for model pretraining.
+Additionally, in accordance with [10], perceptual loss and adversarial loss are also employed for model pretraining.
 III-B Stage 2: Finetuning for Noise Reduction
-Since Stage 1 is trained end-to-end without accounting for channel noise, leading to suboptimal reconstructed results in the presence of noise, we integrate a noise reduction block (NRB) based on a stack of residual Swin Transformer layers, as mentioned in [13] into decoder $G$ to conduct generative image reconstruction.
-Building upon the pretrained model and codebook from Stage 1, as depicted in Fig.
-2 (Left), Stage 2 introduces an physical communication channel module (refer to Section II-B for details) between transmitter and receiver for physical channel simulation.
-Then the noise reduction block is incorporated after the retrieval feature $\mathbf{\hat{z}}_{q}$ to refine semantic features, yielding the refined semantic feature $\mathbf{\hat{z}}_{r}$.
+Since Stage 1 is trained end-to-end without considering any channel noise, leading to suboptimal reconstructed results in the presence of noise, we integrate a noise reduction block (NRB) based on a stack of residual Swin Transformer layers, as mentioned in [11] into decoder $G$ to conduct generative image reconstruction.
+Building upon the pretrained model and multilevel codebook from Stage 1, as depicted in Fig.
+2 (Left), Stage 2 introduces a physical communication channel (refer to Section II-B for details) between transmitter and receiver for physical channel simulation.
+Then the proposed NRB is incorporated after the retrieval noisy feature $\mathbf{\hat{z}}_{q}$ to refine semantic features, yielding the refined semantic feature $\mathbf{\hat{z}}_{r}$.
 Feature Requantization
-To further enhance the quality of $\mathbf{\hat{z}}_{r}$, we propose requantizing it to obtain $\mathbf{\tilde{z}}_{r}$, formulated as follows:
+To further enhance the quality of $\mathbf{\hat{z}}_{r}$, we propose requantizing it to obtain $\mathbf{\tilde{z}}_{r}$ by:
 $\mathbf{\tilde{z}}_{r}=\mathcal{MOC-RVQ}(\mathbf{\hat{z}}_{r}).$ (9)
-The intuition behind this lies in the fact that MOC is trained in the absence of any noise, establishing itself as a high-quality semantic knowledge base to inject high quality prior for better feature restoration.
+The intuition behind this lies in the fact that MOC-RVQ is trained in the absence of any noise, establishing itself as a high-quality semantic knowledge base to inject high quality prior for better feature restoration.
 Training Objectives
-For NRB finetuning, we generate the noiseless quantized semantic feature of image $\mathbf{I}$ through $\mathbf{z}_{q}=\mathcal{MOC}(E(\mathbf{I}))$ as the ground-truth feature.
-This ground-truth feature is then used to calculate the Gram matrix loss and L2 loss for the refined semantic feature $\mathbf{\hat{z}}_{r}$ by the following formula:
+For NRB finetuning, we first generate the noiseless quantized semantic feature of image $\mathbf{I}$ through $\mathbf{z}_{q}=\mathcal{MOC-RVQ}(E(\mathbf{I}))$ as ground-truth feature.
+Then, this ground-truth feature is utilized to calculate the Gram matrix loss and L2 loss for the refined semantic feature $\mathbf{\hat{z}}_{r}$ by:
 $\mathcal{L}_{NR}=\|\psi(\mathbf{\hat{z}}_{r})-\psi(sg[\mathbf{z}_{q}])\|_{2}^{2}+\alpha\|\mathbf{\hat{z}}_{r}-sg[\mathbf{z}_{q}]\|_{2}^{2},$ (10)
@@
 The first term of $\mathcal{L}_{NR}$, also known as style loss, has been demonstrated to be effective for improved feature recovery [11].
-Note that the model and codebook trained in Stage 1 remain fixed, and we only finetune the NRB for noise reduction.
+Moreover, we incorporate image-level $L_{1}$ loss for training supervision.
+It is also noteworthy that the autoencoder and codebook trained in Stage 1 remain fixed, and we only finetune the NRB for noise reduction.
 III-C Codebook Reordering
-As discussed in Section I, a discrepancy exists in the local relationship between code indices and code vectors, rendering VQ-based semantic communication systems susceptible to channel noise.
-In response to this challenge, we propose a heuristic solution through the codebook reordering (CR) algorithm outlined in Algorithm 1.
-The algorithm systematically searches for the nearest neighbor vector within $\mathcal{C}$, constructing a new vector sequence, which is then assembled to form a preliminarily sorted codebook.
-Furthermore, to align with the design principles of Gray code mapping in digital constellation modulation, we also integrate Gray code mapping into our codebook reordering algorithm.
+As discussed in Section I, a discrepancy exists in the local relationship between code indices and code vectors.
+In response to this challenge, before commencing Stage 2, we propose a heuristic solution through a codebook reordering (CR) algorithm outlined in Algorithm 1.
+The algorithm systematically searches for the nearest neighbor vector within $\mathcal{C}$, resulting in a new vector sequence, which is then assembled to form a preliminarily sorted codebook.
+Furthermore, to align with the Gray code mapping process in digital constellation modulation, we also integrate Gray code mapping into our CR algorithm (optional).
 The basic idea is to establish proximity relationships under the context of Gray code mapping.
-For instance, given a 3-bit Gray code sequence represented in decimals as $g=\{0,1,3,2,6,7,5,4\}$, it becomes apparent that "2" and "6" are neighboring elements within the framework of Gray code mapping.
+For instance, given a 3-bit Gray code sequence represented in decimals as $g=\{0,1,3,2,6,7,5,4\}$, it becomes apparent that "2" and "6" are neighboring elements.
 Thus, it is essential to allocate two closely related code vectors to these indices.
+1
 input : A disordered codebook $\mathcal{C}$ of size $N\times n_{q}$
 output : An ordered codebook $\mathcal{C}^{*}$ of size $N\times n_{q}$
+2
 // Initialization
@@
 8       Append $\mathbf{s}_{i}$ to $\mathcal{C}^{*}$ ;
+9
 // Gray Code Mapping
@@
 11 $\mathcal{C}^{*}[g,:]\leftarrow$ $\mathcal{C}^{*}$ ;
-9
 Algorithm 1 Codebook Reordering
 After codebook reordering, the semantic distance between two adjacent Gray code vectors in the updated codebook $\mathcal{C}^{*}$ is reduced compared to the previous disordered version.
-This addresses the issue where a small jump in the code index could lead to a significant jump in the semantic vector.
+This addresses the issue where a small jump in the code index would lead to a significant jump in the semantic vector.
 IV Experiments
+In this section, we empirically demonstrate the qualitative and quantitative results of our model, conduct model ablation, and compare our model with different baselines.
 (a) PSNR (b) SSIM (c) LPIPS
-Figure 3: Experimental comparison of methods using PSNR, SSIM, and LPIPS metrics over an additive white Gaussian noise (AWGN) channel with SNR varying from -5 to 30.
-L1 to L4 denote the level of quantization of our proposed MOC-RVQ, which directly determines the number of bits that need to be transmitted for communication.
+Figure 3: Experimental comparison using PSNR, SSIM, and LPIPS metrics over AWGN channels with SNR from -5 to 30.
+L1 to L4 correspond to different quantization levels of MOC-RVQ, with L1 as a baseline for the VQGAN-based variant.
+JSCC-96 is an analog baseline, with a CBR six times higher than L4, and requires analog modulation or full-resolution constellation, which hinders compatibility with current digital communication systems.
 (a) PSNR (b) SSIM (c) LPIPS
 Figure 4: Ablation study to comprehend the influence of our proposed noise reduction block (NRB) and codebook reordering (CR) algorithm.
-Note that all experiments are conducted on the basis of L4 transmission.
-IV-A Implementation Details
+Note that all experiments are conducted under L4 transmission, and ’w/o’ denotes ’without’.
+There are clear gaps between different setting, which verifies the effectiveness of the proposed NRB and CR designs.
+IV-A Experimental Setups
 Model and Dataset
-In accordance with Chen et al.
-[13], we employ a baseline autoencoder with a downsampling factor $f=8$, indicating that the resolution of latent features is $1/f^{2}$ of the input resolution.
-To ensure effective model training, we assemble a comprehensive training set by integrating diverse datasets: DIV2K [14], Flickr2K [15], DIV8K [16], and a dataset comprising 10,000 facial images from FFHQ [17].
-The generation of training patches involves the following steps: (1) extraction of non-overlapping patches with dimensions $512\times 512$; (2) filtration of patches with minimal texture variation; (3) for the FFHQ dataset, a random resizing operation is applied with scale factors in the range [0.5, 1.0] before cropping.
-For testing, we uniformly sample 25 high resolution images (with average resolution of $2012\times 1407$) from DIV2K validation set as our test set to simulate modern image transmission scenarios.
+Similar to [11], we utilize the autoencoder from VQGAN [10] with a downsampling factor of $f=8$, meaning that the resolution of the bottleneck feature is $1/f^{2}$ of the input, as the backbone.
+For effective model training, we assemble a comprehensive training set by integrating: DIV2K [12], Flickr2K [13], DIV8K [14], and a dataset comprising 10,000 facial images from FFHQ [15].
+Then the training patches are generated by the following steps: (1) Extraction of non-overlapping patches with a resolution of $512\times 512$; (2) Filtration of patches with minimal texture variation; (3) For the FFHQ dataset, a random resizing operation is applied with scale factors in the range [0.5, 1.0] before cropping.
+For testing, we evenly sample 25 high-resolution images ($2012\times 1407$ in average) from the DIV2K validation set to simulate modern image transmission scenarios.
 Training Details
-We employ the Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.99$ for both stages, maintaining a learning rate of $0.0001$ throughout the training process.
-The input image size is randomly cropped to $256\times 256$ with a batch size of 16.
-The model pretraining stage requires approximately 3 days on 2 GeForce RTX 3090 GPUs, and the NRB finetuning stage also takes about 3 days on the same device.
+We employ an Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.99$, and maintain a learning rate of $0.0001$ throughout the training process.
+The input image size is randomly cropped to $256\times 256$ for efficient training with a batch size of 16.
+The model pretraining stage requires approximately 3 days on 2 GeForce RTX 3090 GPUs, while the NRB finetuning stage also takes about 3 days on the same device.
+Channel Settings and Evaluation Metrics
+We simulate AWGN channels with SNR varying from -5 to 30 to encompass low, medium, and high-quality channel conditions.
+Note that the proposed NRB is finetuned under these AWGN channels and all digital models are equipped 64-QAM modulation.
+For evaluation, PSNR, SSIM, and LPIPS are utilized as metrics for various methods.
+LPIPS [16] is specifically introduced for evaluating the perceptual quality of generated images under the semantic communication scenarios.
 IV-B Simulations and Discussions
-Figure 5: Visualizing the impact of NRB across varying channel conditions.
-Even in poor channel quality (SNR=5), our full model impressively reconstructs images with clear semantic meaning, demonstrating the robustness of VQ-based semantic communication systems with the assistance of NRB.
-In this section, we empirically demonstrate the qualitative and quantitative results of our approach, conduct model ablations, and compare our method with two state-of-the-art baselines.
-Note that all approaches are under 64-QAM.
-For channel simulation, initially, we simulate an additive white Gaussian noise (AWGN) channel with SNR varying from -5 to 30 to encompass low, medium, and high-quality channel conditions.
-Subsequently, we finetune our proposed NRB under this AWGN channel to enhance image reconstruction.
-For comparison, we apply the better portable graphics (BPG [18]) and the joint picture expert group (JPEG [19]) as baselines.
+Figure 5: Visualization of removing both the NRB and CR across varying channel conditions.
+Even in poor channel quality (SNR=0), our full model impressively reconstruct images with clear meaning, demonstrating the robustness of VQ-based semantic communication systems with the assistance of NRB and CR.
+We apply the better portable graphics (BPG [17]) and the joint picture expert group (JPEG [18]) as digital baselines by setting the number of bits after compression similar to our proposed MOC-RVQ.
 Notably, both of these image codec formats are susceptible to bit errors, leading to potential image format errors and decoding failures.
 To mitigate this, we introduce a widely used channel error correction code called Low-Density Parity-Check (LDPC) by setting the LDPC block length and bit rate as 648 bits and 1/2, respectively.
-Thanks to the self-restoring capability of our proposed method, we do not apply any channel error correction code, which reduces the number of transmitted bits and significantly improves the transmission efficiency of the proposed communication system.
-As depicted in Fig.
-3, we utilize PSNR, SSIM, and LPIPS as metrics for various methods with a similar number of transmitted symbols over the AWGN channel.
-LPIPS [20] is specifically introduced for evaluating the perceptual quality of generated images in semantic communication scenarios.
-Our proposed method exhibits remarkable robustness in deteriorating channel quality, outperforming traditional image coding even without channel error correction mechanism.
-This is attributed to the VQ-based semantic communication system’s independence from encoding formats, where minor bit changes do not lead to decoding failures.
-Additionally, the incorporation of NRB and codebook reordering further enhances the robustness of the semantic communication system.
-Figure 3 also reveals that lowering the transmission level (L4 to L1) of our model reduces the number of transmitted bits, but it also compromises reconstruction performance.
-Especially, we can clearly observe the gaps at different levels in LPIPS metrics, which implies that we can dynamically adjust the number of bits that need to be transmitted based on the quality of the channel and the receiver’s requirements for image quality.
-In Figure 4, we present an ablation study to understand the impact of our proposed NRB and the CR algorithm.
-Focusing on LPIPS scores, which better capture perceptual quality, we observe the crucial role of NRB in reconstruction performance, especially under poor channel quality.
-Simultaneously, the proposed codebook reordering algorithm provides additional gains in image reconstruction performance.
-Figure 5 illustrates the visual effect of NRB.
-With the assistance of NRB, even under relatively poor channel quality (SNR=5), our full model impressively reconstructs images with clear semantic meaning, which greatly expands the working conditions of VQ-based semantic communication systems.
+Thanks to the self-restoring capability of our proposed method, we do not apply any channel error correction code for MOR-RVQ.
+Additionally, we implement an analog JSCC baseline, referred to as JSCC-96, by retaining only the autoencoder from MOC-RVQ and removing all its digital components.
+Without loss of generality, we set the number of channels in the bottleneck feature (the encoder output) to 96 for comparison, resulting in a channel bandwidth ratio (CBR) that is six times greater than that of MOC-RVQ-L4.
+The definition of CBR is as follows:
+$CBR=\frac{B}{H\times W\times 3},$ (11)
+where $B$ denotes the number of transmitted complex symbols, and $H,W$ represent the width and height of the input RGB image, respectively.
+Note that the symbols of analog scheme require full-resolution constellation, while the symbols of our method only need few discrete states.
+From Fig.
+3, we can observe that our proposed method exhibits robustness in deteriorating channel quality when compared with traditional image coding, even in the absence of a channel error correction mechanism.
+Additionally, while the analog scheme JSCC-96 demonstrates commendable performance, it requires a higher CBR and faces compatibility challenges with existing digital communication systems.
+Figure 3 also shows that reducing the transmission level (from L4 to L1) in our model, which directly decreases the number of transmitted bits, leads to a compromise in reconstruction performance.
+Notably, the gaps at different levels are clearly visible in the LPIPS metrics.
+It is also worth mentioning that L1 can serve as a special VQGAN-based digital baseline for comparison.
+Additionally, We also conduct ablation study (illustrated in Fig.
+4) to assess the effects of our proposed NRB and the CR algorithm.
+We can observe that the proposed NRB plays a crucial role in the reconstruction performance.
+Simultaneously, the proposed CR algorithm contributes additional gains to the image reconstruction performance.
+Figure 5 illustrates the visual effect of removing both the NRB and CR.
+With the assistance of both these designs, even under relatively poor channel quality (SNR=0), our full model can impressively reconstruct images with clear semantic meaning, thus significantly expanding the working conditions of VQ-based semantic communication systems.
 V Conclusion
-We introduce MOC-RVQ, a novel two-stage framework addressing challenges in vector quantization and digital constellation modulation for semantic communication systems.
-By incorporating a multi-head octonary codebook (MOC) to compress the index range and a residual vector quantization (RVQ) mechanism for efficient multilevel semantic communication, our approach demonstrates superior performance over existing methods like BPG and JPEG, even without channel error correction coding.
-The innovative two-stage model training framework, featuring a noise reduction block (NRB), enhances the system’s ability to restore high-quality semantic features in the presence of channel noise.
-Additionally, a heuristic codebook reordering algorithm improves system robustness against channel noise, mitigating the inherent mismatch between vector quantization and digital constellation modulation.
-MOC-RVQ represents a significant advancement in generative semantic communication, providing a promising solution for the joint design challenges and paving the way for further exploration in intelligent communication systems.
+In this work, we propose MOC-RVQ, a novel approach that combines a multi-head octonary codebook (MOC) with residual vector quantization (RVQ) in a two-stage training framework to address the challenges between vector quantization and digital constellation modulation.
+By leveraging MOC for codebook compression and RVQ for efficient multilevel semantic communication, our method surpasses traditional techniques like BPG and JPEG, and achieves performance comparable to analog JSCC.
+Furthermore, the two-stage framework, enhanced by the NRB design and CR algorithm, significantly improves the system’s ability to recover high-quality semantic features under channel noise.
+While effective, MOC-RVQ is specifically optimized for 64-QAM modulation, suggesting the need for codebook adjustments and retraining for other modulation schemes, providing avenues for future research.